OpenStudy (anonymous):

Bette is planning to drive 567 miles on a trip to a national park. If she plans to average a speed of 68 mph, how many hours will it take her to get to the park? Round your answer to the nearest tenth of an hour.

OpenStudy (anonymous):

Time = distance/speed (T = D/S) T = 567/68 Now just divide 567 by 68 and you should get you hours :)